OpenStudy (anonymous):

guy wire How long is a guy wire reaching from the top of a 3ft pole to a point on the ground 2 feet from the pole?

OpenStudy (anonymous):

use the Pythagorean theorem here. c^2=a^2+b^2

OpenStudy (anonymous):

in this case a=4 and b=2, thus: c^2=4^2+2^2 c^2=16+4 c^2=20 take the sqrt of both sides and we get: c=sqr20 further evluating: c=2sqrt5

OpenStudy (anonymous):

thus the wire is about 4.472 feet long or extacly 2sqrt5 feet long